# Science Time Series: Deep Learning in Hydrology

### v1

@@ Abstract
This research is part of a systematic study of scientific time series. In the last three years, hundreds of papers and over fifty new deep-learning models have been described for time series models. These mainly focus on the key aspect of time dependence, whereas in some scientific time series, the situation is more complex with multiple locations, each location having multiple observed and target time-dependent streams and multiple exogenous (known) properties that are either constant or time-dependent. Here, we analyze the hydrology time series using the CAMELS and Caravan global datasets on catchment rainfall and runoff. Together, these have up to 6 observed streams and up to 209 static parameters defined at each of about 8000 locations. This analysis is fully open source with a Jupyter Notebook running on Google Colab for both an LSTM-based analysis and the data engineering preprocessing. Our goal is to investigate the importance of exogenous data, which we look at using eight different choices on representative hydrology tasks. Increasing the exogenous information significantly improves the data representation, with the mean square error decreasing to 60% of its initial value in the largest dataset examined. We present the initial results of studies of other deep-learning neural network architectures where the approaches that can use the full observed and exogenous observations outperform less flexible methods, including Foundation models. Using the natural annual periodic exogenous time series produces the largest impact, but the static and other periodic exogenous streams are also important. Our analysis is intended to be valuable as an educational resource and benchmark.
I Introduction
I-A Spatio-temporal Series Datasets
@@
Scientific data is frequently represented as spatio-temporal series, where time series data are often influenced by geographical factors. The language of spatio-temporal series is used as a common application type, where the ”series” can refer to any ordered sequential data points. These sequences can belong to any collection (bag), not restricted to Euclidean space-time, as long as sequences are labeled in some way and have properties that are consequent to the label. In the case of COVID-19 data [1, 2], daily case / death statistics are grouped by location (e.g. city, county, country) and influenced by demographic characteristics of these locations [3]. In the case of earthquake data [4], the earthquakes are grouped by 11 km x 11 km regions [5]. Similarly, in the case of hydrology data, daily precipitation and streamflow are grouped by catchments and are affected by environmental attributes and locations of these sites.
@@
Typically, the data in these contexts are recorded as space-time-stamped events. Fig 1. However, data can be converted into spatio-temporal series by binning in space and time. Comparing deep learning for time series with coupled ordinary differential equations for multi-particle systems motivates the use of an evolution operator to describe the time dependence of complex systems. Our research views deep learning applied to spatio-temporal series as a method for identifying the time evolution operator governing the behavior of complex systems. Metaphorically, the training process uncovers hidden variables representing the system’s underlying theory, similar to Newton’s laws. Previous studies on COVID-19 [3] and earthquakes [5] show neural networks’ ability to learn spatiotemporal dependencies in spatio-temporal series data. This work extends this approach to hydrology, demonstrating deep learning’s ability to model the rainfall-runoff process.
@@
I-B Rainfall-Runoff Problem
Rainfall-runoff modeling, a key challenge in hydrology, aims to model the physical process by which water on land surface (precipitation or snowmelt) moves to streams [6]. As part of the hydrologic cycle, precipitation on land either evaporates, transpires, infiltrates to recharge groundwater, or becomes surface runoff entering a catchment [6]. A catchment, or watershed, is an area where all precipitation collects and drains into a common outlet, such as a river, lake, or reservoir [6]. Surface runoff contributes to streamflow, which is defined as the volumetric discharge that takes place in a stream or channel. Runoff $RO$ can be estimated with Equation. 1, where $Q$ is streamflow, and $GW_{out}$ is ground-water outflow. Stream outflow $Q$ can be estimated with Equation. 2, where $P$ is precipitation, $GW_{in}$ is ground-water inflow, $ET$ is evapotranspiration, $\Delta S$ is the change in liquid and solid forms of storage. Eventually, this streamflow and ground water outflow reaches the ocean, where it evaporates, condenses into clouds, and returns as rainfall on land, completing the hydrologic cycle [6].
@@ I-B Rainfall-Runoff Problem
Rainfall, like many natural phenomena, is periodic, with daily patterns that follow consistent seasonal cycles. While streamflow result from precipitation, it is also influenced by static environmental factors such as soil type, land cover, slope, etc. Based on this, we hypothesize that daily streamflow in a catchment can be predicted using a combination of daily meteorological forcing data, and the spatially and temporally distributed hydrologic, climatologic, geologic, pedologic, and land-use data [6]. Additionally, a neural network can learn the seasonal patterns of these hydrological processes to produce accurate forecasts.
I-C Related Work
@@
Traditional rainfall-runoff modeling has typically focused on individual catchments. The first documented model, introduced in 1851, used linear regression to predict discharge from precipitation intensity and runoff [7]. Since then, scientific advancements have led to more sophisticated models based on mathematical formulas and physical laws. The advent of computers brought digital hydrological models. At a time when computers are expensive, slow, and low in memory, the Stanford Watershed Model [8] was proposed. It was seen as one of the first and most successful digital computer models [9]. As computers become more powerful, distributed models [10, 11] emerged, allowing for hydrological models to closely couple to geographical information systems for the input data [9]. Taking advantage of the number of parameters offered, these physical-based distributed models perform exceptionally well. However, the high computational cost to calibrate these parameters, and the limited availability of data hinder their use in large-scale forecasting applications [12].
@@
Groundbreaking advancements in deep learning models and the publication of structured large-sample Hydrology dataset have overcome this limitation, enabling the study of nation or global scale rainfall-runoff modeling [13, 14]. In the 1990s, Artificial Neural Network (ANN) based rainfall-runoff model was proposed [15, 16, 17]. Although scientists were initially hesitant to embrace this novel “black box” approach due to the lack of extensive studies, deep learning based Hydrology models proved successful [17]. Its exceptional capability in simulating complex non-linear systems is particularly advantageous for Hydrology modeling. In 2018, the focus of the field shifted towards Long Short-Term Memory (LSTM) based models, which excelled in learning sequential dependencies within time series data [13] [18]. These models have shown great success in large-scale hydrological time series predictions. Further studies explored the interpretability of such LSTM models within physics context [19]. Since then, an open source library for the LSTM based rainfall-runoff model was published [20].
II Data And Methods
II-A Datasets selection
@@
Recent deep learning studies on Hydrology have been driven by the advent of CAMELS (Catchment Attributes and Meteorology for Large-sample Studies) datasets, which established a standard for organizing big Hydrology data at different nations across the globe. The first CAMELS dataset [21], initially proposed with 671 catchments in the U.S., benchmarked the types of static and time series properties necessary for large sample hydrology datasets containing hundreds of catchments or more. The high dimensionality of static data along with the 20-year duration of daily time series data made it suitable for nation-scale Hydrology modeling using neural network models. Since then, CAMELS-standard datasets have been published for countries including the United Kingdom [22], Chile [23], Australia [24], Brazil [25], Switzerland [26], Sweden [27], France [28], etc.
II-B Three-Nation Combined CAMELS Data
Similarities shared by CAMELS-standard national datasets allow for the combination of national datasets into a large global dataset, which can be used for global-scale training. In this study, the experimental dataset was produced by combining CAMELS data from three nations: the United States[21], United Kingdom[22], and Chile. These datasets were selected as they provide the earliest available CAMELS-structured data during the data preprocessing phase of this study. We select static features and time series targets that are shared across the three datasets prior to combination.
TABLE I: Three-Nation Combined CAMELS Data Source
CAMELS-US CAMELS-GB CAMELS-CL
Forcing a Maurer [29] CEH-GEAR [30] CR2MET [31]
CHESS-met [32]
Streamflow USGS [33] NRFA [34] CR2 [35]
@@ II-B Three-Nation Combined CAMELS Data
II-C CAMELS Preprocessing
For dynamic variables, we selected the common interval of 7,031 days, spanning from October 2, 1989, to December 31, 2008. The start date of a water year, as defined by the U.S. Geological Survey, is October 1st [6]. To account for time differences between the nations, we shifted the training data by one day, beginning on October 2nd. An analysis of NaNs in data revealed no NaN values in time series data, yet some exists in the static features. Since missing data constitutes only a small percentage of total static data and occurs miscellaneously, it is filled with the mean value of that attribute from all catchments. Categorical variables are only present in the form of months from January to December. We encode them with ordinal encoding between 0 and 1 to preserve the natural order of months.
II-D Caravan
Published in 2023, the Caravan dataset [36], consists of seven preprocessed CAMELS-standard national datasets that contain identical static exogenous features and time series properties. Caravan aggregates data from 6,830 catchments across 16 nations spanning four continents, making it ideal for global rainfall-runoff modeling with large-scale hydrological data.
TABLE II: Caravan Sub-datasets
@@ II-D Caravan
CAMELS (US) [21] 482
CAMELS-AUS [24] 150
CAMELS-BR [25] 376
CAMELS-CL [23] 314
CAMELS-GB [22] 408
HYSETS (North America) [37] 4621
LamaH-CE [38] 479
The Caravan dataset is not constructed by identifying common static and time series properties shared by previously published versions of each sub-dataset. Instead, it is derived from data sources that differ significantly from those used in the original CAMELS datasets for various nations. While the catchments included in the Caravan sub-datasets are the same as those in the original versions, the actual data is sourced from global hydrological sources rather than the local sources used in the original publications. This shift to global sources enables the collection of more standardized and abundant data, making it suitable for global-scale comparative hydrological studies—one of the key motivations behind the creation of Caravan. However, it is important to note that data from global sources may vary slightly from local sources, which are often more precise. For example, in the CAMELS-US sub-dataset within Caravan, all forcing data is sourced from ERA5-Land [39], streamflow data from GSIM [40] [41], and static properties from HydroATLAS [42]. In contrast, the original CAMELS-US [21] dataset included static and time series data from three sources: NLDAS [43], Maurer [29], and DayMet [44], with streamflow data from USGS [33]. After comparison, only the streamflow data remains consistent between CAMELS-US in the original publication and in Caravan. The correlation coefficient for precipitation time series data between different data sources is presented in III. Nevertheless, the Caravan paper [36] addresses this concern, noting that the correlations between Caravan and each of the three CAMELS-US data products are not consistently lower than the correlations within the individual CAMELS-US data products.
TABLE III: CAMELS-US Caravan Comparison
@@
III Methods
III-A Long Short-Term Memory
The Long Short-Term Memory (LSTM) network [45] is a specialized variant of a Recurrent Neural Network (RNN) designed to address the vanishing gradient problem through its unique memory cell structure. In an LSTM block (as shown in the figure below), the cell state (denoted by $C$) serves as long-term memory. Minimal weight updates to the cell state during backpropagation effectively mitigate the vanishing gradient problem, making LSTM particularly well-suited for time series tasks involving long sequences of data. Additionally, the hidden state (denoted by $h$) serves as short-term memory. Weights (denoted by $W$) and biases (denoted by $b$) are applied to the input and passed through sigmoid and tanh activation functions. The output from the hidden state is then used to update the cell state, while the updated cell state, in turn, informs the hidden state during the final stage of computation within the LSTM block.
@@ III-A Long Short-Term Memory
$f_{t}=sigmoid(W_{f}\times x_{t}+U_{f}\times h_{t-1}+b_{f})\vspace{-5mm}$ (3)
$i_{t}=sigmoid(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vspace{-5mm}$ (4)
$\hat{C}_{t}=tanh(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vspace{-5mm}$ (5)
$O_{t}=sigmoid(W_{o}\times x_{t}+U_{o}\times h_{t-1}+b_{o})\vspace{-5mm}$ (6)
$C_{t}=C_{t-1}\times f_{t}+i_{t}\times\hat{C}_{t}\vspace{-5mm}$ (7)
@@
III-B Model Setup
The LSTM model used in this study is implemented using the Tensorflow framework, which allows for customization of layer parameters. The baseline model architecture is shown in Fig. 3. Detailed activation function setup is shown in Tab. IV. To prevent overfitting, a dropout rate of 20% is applied to the layers.

### v2

# Deep Learning Foundation and Pattern Models: Challenges in Hydrological Time Series

@@ Abstract
There has been active investigation into deep learning approaches for time series analysis, including foundation models. However, most studies do not address significant scientific applications. This paper aims to identify key features in time series by examining complex hydrology data. Our work advances computer science by emphasizing critical application features and contributes to hydrology and other scientific fields by identifying modeling approaches that effectively capture these features. Scientific time series data are inherently complex, involving observations from multiple locations, each with various time-dependent data streams and exogenous factors that may be static or time-varying and either application-dependent or purely mathematical. This research analyzes hydrology time series from the CAMELS and Caravan global datasets, which encompass rainfall and runoff data across catchments, featuring up to six observed streams and 209 static parameters across approximately 8,000 locations. Our investigation assesses the impact of exogenous data through eight different model configurations for key hydrology tasks. Results demonstrate that integrating exogenous information enhances data representation, reducing mean squared error by up to 40% in the largest dataset. Additionally, we present a detailed performance comparison of over 20 state-of-the-art pattern and foundation models. The analysis is fully open-source, facilitated by Jupyter Notebook on Google Colab for LSTM-based modeling, data preprocessing, and model comparisons. Preliminary findings using alternative deep learning architectures reveal that models incorporating comprehensive observed and exogenous data outperform more limited approaches, including foundation models. Notably, natural annual periodic exogenous time series contribute the most significant improvements, though static and other periodic factors are also valuable. This research serves as both an educational tool and benchmark resource.
keywords:
Deep Learning, Hydrology, Time Series Analysis, Long Short-Term Memory (LSTM), Foundation Models
1 Introduction
1.1 Spatio-temporal Series Datasets
@@
Scientific data is frequently represented as spatio-temporal series, where time series data are often influenced by geographical factors. The language of spatio-temporal series is used as a common application type, where the ”series” can refer to any ordered sequential data points. These sequences can belong to any collection (bag), not restricted to Euclidean space-time, as long as sequences are labeled in some way and have properties that are consequent to the label. In the case of COVID-19 data CDC (n.d.); cdc (n.d.), daily case / death statistics are grouped by location (e.g. city, county, country) and influenced by demographic characteristics of these locations Fox et al. (2021). In the case of earthquake data usg (n.d.), the earthquakes are grouped by 11 km x 11 km regions Fox et al. (2022). Similarly, in the case of hydrology data, daily precipitation and streamflow are grouped by catchments and are affected by environmental attributes and locations of these sites.
The spatio-temporal nature of scientific time series data means it is significantly influenced by the spatial variability of environmental factors. In recent years, computer science has seen the publication of hundreds of papers and more than fifty new deep learning models for time series analysis sci (2023); nix (n.d.); ddz (n.d.); Jin et al. (2023); Liang et al. (2024); Jafari et al. (2024), primarily focused on temporal dependencies. However, many of these models overlook the importance of incorporating application-specific environmental or exogenous data. This work addresses these gaps by emphasizing key considerations for developing time series models tailored to data-driven scientific applications.
@@
Typically, the data in these contexts are recorded as space-time-stamped events. Fig 1. However, data can be converted into spatio-temporal series by binning in space and time. Comparing deep learning for time series with coupled ordinary differential equations for multi-particle systems motivates the use of an evolution operator to describe the time dependence of complex systems. Our research views deep learning applied to spatio-temporal series as a method for identifying the time evolution operator governing the behavior of complex systems. Metaphorically, the training process uncovers hidden variables representing the system’s underlying theory, similar to Newton’s laws. Previous studies on COVID-19 Fox et al. (2021) and earthquakes Fox et al. (2022) show neural networks’ ability to learn spatiotemporal dependencies in spatio-temporal series data. This work extends this approach to hydrology, demonstrating deep learning’s ability to model the rainfall-runoff process.
@@
1.2 Rainfall-Runoff Problem
Rainfall-runoff modeling, a key challenge in hydrology, aims to model the physical process by which water on land surface (precipitation or snowmelt) moves to streams Dingman (2015). As part of the hydrologic cycle, precipitation on land either evaporates, transpires, infiltrates to recharge groundwater, or becomes surface runoff entering a catchment Dingman (2015). A catchment, or watershed, is an area where all precipitation collects and drains into a common outlet, such as a river, lake, or reservoir Dingman (2015). Surface runoff contributes to streamflow, which is defined as the volumetric discharge that takes place in a stream or channel. Runoff $RO$ can be estimated with Equation. 1, where $Q$ is streamflow, and $GW_{out}$ is ground-water outflow. Stream outflow $Q$ can be estimated with Equation. 2, where $P$ is precipitation, $GW_{in}$ is ground-water inflow, $ET$ is evapotranspiration, $\Delta S$ is the change in liquid and solid forms of storage. Eventually, this streamflow and ground water outflow reaches the ocean, where it evaporates, condenses into clouds, and returns as rainfall on land, completing the hydrologic cycle Dingman (2015).
@@ 1.2 Rainfall-Runoff Problem
Rainfall, like many natural phenomena, is periodic, with daily patterns that follow consistent seasonal cycles. While streamflow result from precipitation, it is also influenced by static environmental factors such as soil type, land cover, slope, etc. Based on this, we hypothesize that daily streamflow in a catchment can be predicted using a combination of daily meteorological forcing data, and the spatially and temporally distributed hydrologic, climatologic, geologic, pedologic, and land-use data Dingman (2015). Additionally, a neural network can learn the seasonal patterns of these hydrological processes to produce accurate forecasts.
1.3 Related Work
@@
Traditional rainfall-runoff modeling has typically focused on individual catchments. The first documented model, introduced in 1851, used linear regression to predict discharge from precipitation intensity and runoff J. (1851). Since then, scientific advancements have led to more sophisticated models based on mathematical formulas and physical laws. The advent of computers brought digital hydrological models. At a time when computers are expensive, slow, and low in memory, the Stanford Watershed Model Crawford (1966) was proposed. It was seen as one of the first and most successful digital computer models Beven et al. (2012). As computers become more powerful, distributed models Abbott et al. (1986); Moore and Clarke (1981) emerged, allowing for hydrological models to closely couple to geographical information systems for the input data Beven et al. (2012). Taking advantage of the number of parameters offered, these physical-based distributed models perform exceptionally well. However, the high computational cost to calibrate these parameters, and the limited availability of data hinder their use in large-scale forecasting applications Sitterson et al. (2017).
@@
Groundbreaking advancements in deep learning models and the publication of structured large-sample Hydrology dataset have overcome this limitation, enabling the study of nation or global scale rainfall-runoff modeling Kratzert et al. (2018); Shen (2018). In the 1990s, Artificial Neural Network (ANN) based rainfall-runoff model was proposed Hsu et al. (1995); Tokar A. Sezin and Johnson Peggy A. (1999); Govindaraju and Rao (2013). Although scientists were initially hesitant to embrace this novel “black box” approach due to the lack of extensive studies, deep learning based Hydrology models proved successful Govindaraju and Rao (2013). Its exceptional capability in simulating complex non-linear systems is particularly advantageous for Hydrology modeling. In 2018, the focus of the field shifted towards Long Short-Term Memory (LSTM) based models, which excelled in learning sequential dependencies within time series data Kratzert et al. (2018) Hu et al. (2018). These models have shown great success in large-scale hydrological time series predictions. Further studies explored the interpretability of such LSTM models within physics context Kratzert et al. (2019). Since then, an open source library for the LSTM based rainfall-runoff model was published Kratzert et al. (2022).
2 Data And Methods
2.1 Datasets selection
@@
Recent deep learning studies on Hydrology have been driven by the advent of CAMELS (Catchment Attributes and Meteorology for Large-sample Studies) datasets, which established a standard for organizing big Hydrology data at different nations across the globe. The first CAMELS dataset Addor et al. (2017), initially proposed with 671 catchments in the U.S., benchmarked the types of static and time series properties necessary for large sample hydrology datasets containing hundreds of catchments or more. The high dimensionality of static data along with the 20-year duration of daily time series data made it suitable for nation-scale Hydrology modeling using neural network models. Since then, CAMELS-standard datasets have been published for countries including the United Kingdom Coxon et al. (2020), Chile Alvarez-Garreton et al. (2018), Australia Fowler et al. (2021), Brazil Chagas et al. (2020), Switzerland Höge et al. (2023), Sweden Teutschbein (2024), France Delaigue et al. (2022), etc.
2.2 Three-Nation Combined CAMELS Data
Similarities shared by CAMELS-standard national datasets allow for the combination of national datasets into a large global dataset, which can be used for global-scale training. In this study, the experimental dataset was produced by combining CAMELS data from three nations: the United StatesAddor et al. (2017), United KingdomCoxon et al. (2020), and Chile. These datasets were selected as they provide the earliest available CAMELS-structured data during the data preprocessing phase of this study. We select static features and time series targets that are shared across the three datasets prior to combination.
Table 1: Three-Nation Combined CAMELS Data Source
Forcing a Streamflow
CAMELS-US Maurer Livneh et al. (2013) USGS Lins (2012)
CAMELS-GB CEH-GEAR Keller et al. (2015) NRFA Hannaford (2004)
CHESS-met Robinson et al. (2017)
CAMELS-CL CR2MET Boisier (2023) CR2 for Climate and Research
@@ 2.2 Three-Nation Combined CAMELS Data
2.3 CAMELS Preprocessing
For dynamic variables, we selected the common interval of 7,031 days, spanning from October 2, 1989, to December 31, 2008. The start date of a water year, as defined by the U.S. Geological Survey, is October 1st Dingman (2015). To account for time differences between the nations, we shifted the training data by one day, beginning on October 2nd. An analysis of NaNs in data revealed no NaN values in time series data, yet some exists in the static features. Since missing data constitutes only a small percentage of total static data and occurs miscellaneously, it is filled with the mean value of that attribute from all catchments. Categorical variables are only present in the form of months from January to December. We encode them with ordinal encoding between 0 and 1 to preserve the natural order of months.
2.4 Caravan
Published in 2023, the Caravan dataset Kratzert et al. (2023), consists of seven preprocessed CAMELS-standard national datasets that contain identical static exogenous features and time series properties. Caravan aggregates data from 6,830 catchments across 16 nations spanning four continents, making it ideal for global rainfall-runoff modeling with large-scale hydrological data.
Table 2: Caravan Sub-datasets
@@ 2.4 Caravan
CAMELS (US) Addor et al. (2017) 482
CAMELS-AUS Fowler et al. (2021) 150
CAMELS-BR Chagas et al. (2020) 376
CAMELS-CL Alvarez-Garreton et al. (2018) 314
CAMELS-GB Coxon et al. (2020) 408
HYSETS (North America) Arsenault et al. (2020) 4621
LamaH-CE Klingler et al. (2021) 479
The Caravan dataset is not constructed by identifying common static and time series properties shared by previously published versions of each sub-dataset. Instead, it is derived from data sources that differ significantly from those used in the original CAMELS datasets for various nations. While the catchments included in the Caravan sub-datasets are the same as those in the original versions, the actual data is sourced from global hydrological sources rather than the local sources used in the original publications. This shift to global sources enables the collection of more standardized and abundant data, making it suitable for global-scale comparative hydrological studies—one of the key motivations behind the creation of Caravan. However, it is important to note that data from global sources may vary slightly from local sources, which are often more precise. For example, in the CAMELS-US sub-dataset within Caravan, all forcing data is sourced from ERA5-Land Muñoz Sabater et al. (2021), streamflow data from GSIM Do et al. (2018) Gudmundsson et al. (2018), and static properties from HydroATLAS Linke et al. (2019). In contrast, the original CAMELS-US Addor et al. (2017) dataset included static and time series data from three sources: NLDAS Xia et al. (2012), Maurer Livneh et al. (2013), and DayMet THORNTON et al. (2014), with streamflow data from USGS Lins (2012). After comparison, only the streamflow data remains consistent between CAMELS-US in the original publication and in Caravan. The correlation coefficient for precipitation time series data between different data sources is presented in 3. Nevertheless, the Caravan paper Kratzert et al. (2023) addresses this concern, noting that the correlations between Caravan and each of the three CAMELS-US data products are not consistently lower than the correlations within the individual CAMELS-US data products.
Table 3: CAMELS-US Caravan Comparison
@@
3 Methods
3.1 Long Short-Term Memory
The Long Short-Term Memory (LSTM) network Hochreiter and Schmidhuber (1997) is a specialized variant of a Recurrent Neural Network (RNN) designed to address the vanishing gradient problem through its unique memory cell structure. In an LSTM block (as shown in the figure below), the cell state (denoted by $C$) serves as long-term memory. Minimal weight updates to the cell state during backpropagation effectively mitigate the vanishing gradient problem, making LSTM particularly well-suited for time series tasks involving long sequences of data. Additionally, the hidden state (denoted by $h$) serves as short-term memory. Weights (denoted by $W$) and biases (denoted by $b$) are applied to the input and passed through sigmoid and tanh activation functions. The output from the hidden state is then used to update the cell state, while the updated cell state, in turn, informs the hidden state during the final stage of computation within the LSTM block.
@@ 3.1 Long Short-Term Memory
$f_{t}=sigmoid(W_{f}\times x_{t}+U_{f}\times h_{t-1}+b_{f})\vspace{-2mm}$ (3)
$i_{t}=sigmoid(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vspace{-2mm}$ (4)
$\hat{C}_{t}=tanh(W_{i}\times x_{t}+U_{i}\times h_{t-1}+b_{i})\vspace{-2mm}$ (5)
$O_{t}=sigmoid(W_{o}\times x_{t}+U_{o}\times h_{t-1}+b_{o})\vspace{-2mm}$ (6)
$C_{t}=C_{t-1}\times f_{t}+i_{t}\times\hat{C}_{t}\vspace{-2mm}$ (7)
@@
3.2 Model Setup
The LSTM model used in this study is implemented using the Tensorflow framework, which allows for customization of layer parameters. The baseline model architecture is shown in Fig. 3. Detailed activation function setup is shown in Tab. 4. To prevent overfitting, a dropout rate of 20% is applied to the layers.